These were accepted.
blog and webpages written. To be kept on action list until DES changes complete.
Done. Craig (and others) have sent lists of possible sites. All have now been added to our 'permitted to scan' list, as have all Informatics subnets. Alison/Ian will continue to work on streamlining process of reporting areas of concern to appropriate individuals.
Action changed - Alison to sort out support-team and rt-team lists. For lists that are to remain, list owners should ensure membership is correct.
Alastair has sorted the problem but it was suggested that the cron job should be run hourly. It will only pick up changes if a machine has been up in the last hour so changes can be missed due to sleep.
Neil is expecting Miles to send a list of requirements that he has gathered. Neil will review and comment/add to as required.
Chris has linked College security pages to our documentation. Chris/Alastair to discuss whether this may cause confusion e.g. conflicting information between College pages and our own.
Stephen has created a script to make it easier to shutdown low criticality machines. It just needs to be tailored to work on individual machine rooms.
To be deferred
Angus responded with further questions. Alison has replied.
MPU will prepare a list of instructions on how to change a DICE machine to being self-managed.
Blog articles suggested:
Blog articles completed:
Report from Computing Executive Group
Reports from units
It was agreed that Toby will send out a reminder to students who have not yet set up their account approximately one week before the start of semester. This year, accounts have started to trickle through from Euclid a little earlier than expected. Next year, we will defer sending the new account details via email until roughly 2 weeks before semester starts.
MPU advised of 'urgent' firmware upgrade for certain PowerEdge models. Check ordershost for details.
The suggestion was made that it might be useful if all Units link their meeting minutes in their reports.
Neil noted that the weekly emails advising of criticality status, profile issues etc have not been received since 12th August. Chris will fix.
It was suggested that Craig should try safari on Windows. Neil confirmed that the documentation has been updated with details of the problem.
Topics for discussion
Things seem to be going to schedule. Stephen explained the changes that he's been making to mod_waklog. Toby will put the version of afs that he has built (1.6.5 for MacOSX 10.5 Leopard) in an accessible place. CSOs will run updaterpms on machines that have been switched off - not CO machines though!
This was trigerred by the situation with mcgahey. MPU will compile a list of steps to take when a DICE machine is to be converted to self-managed.
Discussion led to the conclusion that the edge filter rules which block outbound SMB should remain. Access to Informatics shared space can be through openVPN or the University VPN for IS shares although it was noted that it wouldn't be possible to do both at the same time. A possible option might be to allow users access to rdesktop. There are not many admin users who work from home using Windows laptops so we felt that leaving things as they stand was the best option.
Wednesday 11th September, Turing Room.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh